OpenStudy (anonymous):

The average distance from the earth to the sun is approximately 93,000,000 miles. Write this distance in scientific notation.

Parth (parthkohli):

You start with \(93,000,000 \times 1\). You then shift decimals.. \(\Rightarrow 93 \times 1,000,000\) \(\Rightarrow 9.3 \times 10,000,000\) \(\Rightarrow 9.3 \times 10^{7} \)